# NAPER: Fault Protection for Real-Time Resource-Constrained Deep Neural Networks

### v1

@@ II Related Works
Approach Acc. Avail. Recovery Deployable
Redundancy-based [15, 17, 16, 18, 19] -
@@
Robust DNN [25, 23, 24, 26, 22, 27] - -
Hardware Support [31, 32, 33] -
ECC-based [28, 29, 30] - -
@@ IV-A Model Ensemble Design
NAPER adopts heterogeneous redundancy models that have the same architecture but different parameter values, When combined as an ensemble, these models provide richer data representation and improved accuracy. The ensemble’s accuracy performance depends on model diversity: the more diverse these models are, the better the performance [21, 34, 23]. The simplest yet effective to gain high accuracy is by employing independent training, training the new model independently from scratch with different random parameter initialization [21, 34]. However, independent training can be challenging when the base model is trained in a high-resource environment. Alternatively, fine-tuning some layers [35] or other ensemble generation schemes [36] can be used. Our initial experiments showed that independent training achieves the highest ensemble accuracy, therefore in this paper, NAPER primarily uses it for evaluation.
@@ V-A Datasets and Baselines
We exclude results of leveraging ensemble in conservative methods, such as TMR, because these methods are not designed for ensemble strategy. Ensembling identical redundancy in TMR will not benefit the model’s accuracy. We also exclude ECC-based protection schemes [29, 28] as they are not designed to handle many bit flips.
@@ V-B Software and Hardware Specification
NAPER can be used as a wrapper for any DNN model made with PyTorch. It was primarily tested on an NVIDIA Jetson Nano, a small computer module for edge AI applications. The Jetson Nano features a quad-core ARM Cortex-A57 CPU @ 1.43 GHz, a 128-core Maxwell GPU, and 4 GB of LPDDR4 RAM.
@@ V-D Experiment Scenarios
Timeliness Evaluation: We conducted two evaluations. First, we simulated a video monitoring system and fed protected models with contiguous data at 4 frames per second (fps) for 50 frames, then injected a high fault rate into model parameters five times randomly. We chose 4 fps to meet TMR WCET. At the second evaluation, we varied the bit flip from low to high fault rate at 4 fps and measured the percentage of deadlines met. We define a model as meeting the deadline when it delivers the output before TMR WCET (ignoring accuracy). We used CIFAR10 and averaged over 10 trials.

### v2

@@ Abstract
\$*\$\$*\$footnotetext: Much of the research was conducted when Santriaji was a postdoctoral researcher in the ARPERS group at UCF.
 ©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ II Related Works
Approach
Acc.
Avail.
Recovery Deployable
Redundancy-based
[15, 17, 16, 18, 19]
@@
Robust DNN
[25, 23, 24, 26, 22, 27]
Hardware Support
[31, 32, 33]
ECC-based
[28, 29, 30]
@@ IV-A Model Ensemble Design
NAPER adopts heterogeneous redundancy models that have the same architecture but different parameter values. When combined as an ensemble, these models provide richer data representation and improved accuracy. The ensemble’s accuracy performance depends on model diversity: the more diverse these models are, the better the performance [21, 34, 23]. The simplest yet effective way to gain high accuracy is by employing independent training, training the new model independently from scratch with different random parameter initialization [21, 34]. Independent training can be challenging when training a single model is computationally intensive. Therefore, training redundancy independently will multiply the resource requirements. Alternatively, fine-tuning or transfer learning [35] or other ensemble generation schemes [36] can be used. However, our initial experiments showed that independent training achieves the highest accuracy; therefore, NAPER primarily uses it for evaluation.
@@ V-A Datasets and Baselines
For a fair comparison, all schemes above were implemented at the layer level: fault detection and recovery are conducted after each layer. We exclude results of leveraging an ensemble in conservative methods, such as TMR, because these methods are not designed for ensemble strategy. Ensembling identical redundancy in TMR will not benefit the model’s accuracy. We also exclude ECC-based protection schemes [29, 28] as they are not designed to handle many bit flips.
@@ V-B Software and Hardware Specification
NAPER [^1] https://github.com/ARPERS/NAPER is a wrapper for DNN models made with PyTorch. It was primarily tested on an NVIDIA Jetson Nano, a small computer module for edge AI applications. The Jetson Nano features a quad-core ARM Cortex-A57 CPU @ 1.43 GHz, a 128-core Maxwell GPU, and 4 GB of LPDDR4 RAM.
@@ V-D Experiment Scenarios
Timeliness Evaluation: We conducted two evaluations. First, we simulated a video monitoring system and fed protected models with contiguous data at 4 frames per second (fps) for 50 frames. Then, we injected a high fault rate into model parameters five times randomly, demonstrating the accumulation of bit flips over time. We chose 4 fps to meet TMR WCET. At the second evaluation, we varied the bit flip from low to high fault rate and measured the percentage of deadlines met. We define a model as meeting the deadline when it delivers the output before TMR WCET (ignoring accuracy). We used CIFAR10, averaged over 10 trials.
@@ VII Conclusion
VIII Acknowledgements
This work was partially supported by the National Science Foundation under grant 2106629, Office of Naval Research under grant N00014-23-1-2136, and the Ministry of Education, Culture, Research, and Technology of the Republic of Indonesia through the Garuda Ace Project, with Project IDs 21-I1 and 24-I4.